Tel 4 12. (162); = (x)4. x = (a) 1203 (b) 1201 (c) 1331 (d) 1302 (e) 1221

The Answer is D but how?
Answer #1

162 to base 8 has to converted to base 4 , follow the steps below -

1) First convert 162 base 8 to decimal ie base 10 like below  -

1 x 8^2 + 6 x 8^1 + 2 x 8^0 = 64 + 48 + 2 = 114 is in base 10

So now we have 162 base 8 in decimal as 114 as base 10 -

now conver 114 base 10 to base 4 like below -

Divisor Number to be divided/Number left after dividing Remainder
4 114
4 28 2
4 7 0
4 1 3

So we are left with 1 number left at the last before dividing , and in remainder we have 3,0,2

so merge both this we will get - 1302 which it the answer D in you option.

pls do give a like , thank you